# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 2. PRELIMINARIES
(d) We say that $X\subseteq M$ is coded (in $\mathcal{M)}$ if there is some $c\in M$ such that $\mathrm{Ext}_{\mathcal{M}}(c)=X.$ $X$ is piecewise coded in $\mathcal{M}$ if $X\cap\mathrm{Ext}_{\mathcal{M}}(m)$ is coded for each $m\in M.$
(e) Suppose $\mathcal{M}\models\mathrm{ZF}$. We say that $A\subseteq M$ is $\mathcal{M}$-amenable if $(\mathcal{M},A)\models\mathrm{ZF}(\mathrm{X})$, here $A$ is the interpretation of $X$.
2.3. Definitions. (“geometric shapes” of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures, and $\mathcal{M}$ a submodel [^4] The notion of a submodel here is as in Model Theory; i.e., $\mathcal{M}\subseteq\mathcal{N}$ means $M\subseteq N$ and the interpretation of each nonlogical symbol of $\mathcal{L}$ in $\mathcal{M}$ is the restriction to $M$ of the corresponding of interpretation in $\mathcal{N}$ (in particular, $\mathrm{Ext}_{\mathcal{M}}(a)\subseteq\mathrm{Ext}_{\mathcal{N}}(a)$ for all $a\in M).$ of $\mathcal{N}$ (written $\mathcal{M}\subseteq\mathcal{N)}.$
(a) $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$ if $M^{\ast}=\bigcup\limits_{a\in M}\mathrm{Ext}_{\mathcal{N}}(a).$
(b) Suppose $a\in M.$ $\mathcal{N}$ fixes $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$, and $\mathcal{N}$ enlarges $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)\subsetneq\mathrm{Ext}_{\mathcal{N}}(a).$
(c) $\mathcal{N}$ end extends $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$), if $\mathcal{N}$ fixes every $a\in M$.  End extensions are also referred to in the literature as transitive extensions, and in the old days as outer extensions.
(d) $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$) if $(i)$ $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $(ii)$ for if $a\in M$, $b\in N$, and $\mathcal{N}\models(b\subseteq a)$, then $b\in M$.
(e) $\mathcal{N}$ is a rank extension of $\mathcal{N}$ (written $\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}$) if $\mathcal{N}$ is an end extension of $\mathcal{M}$, and for all $a\in M$ and all $b\in N\backslash M$, $\mathcal{N}\models\mathrm{rank}(a)\in\mathrm{rank}(b)$. Here we assume that $\mathcal{M}$ and $\mathcal{N}$ are models of a sufficient fragment of $\mathrm{ZF}$ (such as $\mathrm{KP}$) in which the rank function is well-defined. $\mathcal{N}$ is a topped rank extension of $\mathcal{M}$ if $\mathrm{Ord}^{\mathcal{N}}\backslash\mathrm{Ord}^{\mathcal{M}}$ has a least element. [^5] Recently rank extensions are also referred to as top extensions, we will not use this terminology as it leads to the expression “a topped top extension”.
(f) $\mathcal{N}$ is a cofinal extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{N)}$ if for every $b\in N$ there is some $a\in M$ such that $b\in\mathrm{Ext}_{\mathcal{N}}(a)$, i.e., $\mathcal{N}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$.
(g) $\mathcal{N}$ is taller than $\mathcal{M}$ if there is some $b\in M$ such that $M\subseteq\mathrm{Ext}_{\mathcal{N}}(b).$ For models of $\mathrm{ZF}$ (or even $\mathrm{KPR}$) this is equivalent to the existence of an “ordinal” in $\mathrm{Ord}^{\mathcal{N}}$ that exceeds all the “ordinals” in $\mathrm{Ord}^{\mathcal{M}}.$
2.4. Definitions. (“logical shapes” of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, and $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures such that $\mathcal{M}$ is a submodel of $\mathcal{N}$.
@@
(b) $\mathcal{N}$ is a self-extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}$) if $(i)$ and $(ii)$ below hold:
$(i)$ $\mathcal{N}$ is interpretable in $\mathcal{M}$, i.e., if $\mathcal{N}$ is isomorphic to a structure $\mathcal{N}^{\ast}$ whose universe $N^{\ast}$, together with the interpretation $R^{\mathcal{N}^{\ast}}$ of each $R\in\mathcal{L}$, are all $\mathcal{M}$-definable.
$(ii)$ There is an $\mathcal{M}$-definable embedding of $\mathcal{M}$ into $\mathcal{N}$.
$\mathcal{N}$ is an elementary self-extension of $\mathcal{M}$ if in $(ii)$ above, the map $j$ is an elementary embedding.
@@
(c) $\mathcal{N}$ is a conservative extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$) if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-definable.
@@
(d) $\mathcal{N}$ is a faithful extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N)}$ if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-amenable (as in Definition 2.2(e)). [^6] The notion of a faithful extension was first introduced and studied in the context of elementary extensions in [En-3].
2.5. Remark. Suppose $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}_{\mathrm{set}}$-structures with $\mathcal{M}\subseteq\mathcal{N}\mathrm{.}$ It is clear that we have the following chain of implications:
@@
$\left(\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N}\right).$
@@
It is also noteworthy that the most common examples of definable extensions in the context of models of set theory are elementary extensions. However, the Boolean valued approach to forcing provides a wealth of definable extensions that are not elementary. More specifically, if $\mathcal{M}$ is a model of $\mathrm{ZFC}$, $\mathbb{B}$ is a complete Boolean algebra from the point of view of $\mathcal{M}$, and $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$ that is in $\mathcal{M}$, then there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into the model $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ (where $\mathcal{M}^{\mathbb{B}}$ is the $\mathbb{B}$-valued forcing extension of the universe, as calculated in $\mathcal{M}$). However, in typical cases $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ are not elementarily equivalent, e.g., we can always arrange $\mathbb{B}$ so that the truth-value of the continuum hypothesis changes in the transition between the $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$. Furthermore, even though $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ might be elementarily equivalent in some cases, it can be verified that in general, the embedding $j$ is not elementary if $\mathbb{B}$ is a nontrivial notion of forcing.
(a) For each $n\in\omega,$ $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{N}$ iff $\mathcal{M}\preceq_{\Pi_{n}}\mathcal{N}.$
(b) If $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, then $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$; thus $\Delta_{1}$-notions are absolute between $\mathcal{M}$ and $\mathcal{N}$, and $\Sigma_{1}$-formulae with parameters from $M$ are preserved in the passage from $\mathcal{M}$ to $\mathcal{N}$.
(c) Suppose $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, $\mathcal{N}$ fixes $a\in M$, and there is bijection $f$ in $\mathcal{M}$ between $a$ and some $b\in M$. Then $\mathcal{N}$ fixes $b$ as well (by $\Delta_{0}$-elementarity, $f$ remains a bijection between $a$ and $b$ as viewed from $\mathcal{N}$).
(d) Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{KPR}$, $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, and $\alpha\in\mathrm{Ord}^{M}$. Then $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{N}}$; this follows from the fact that the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ as a $\Sigma_{1}^{\mathcal{P}}$-formula in $\mathrm{KPR}$ (using the usual recursive definition of the $\mathrm{V}_{\alpha}$-hierarchy).
(e) If $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, and $\mathcal{M}$ and $\mathcal{N}$ both satisfy the axioms of extensionality and powerset, then $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$ (as noted by Forster and Kaye [FS]). Since the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ can be written as a $\Sigma_{1}^{P}$-formula, it is absolute between models $\mathcal{M}$ and $\mathcal{N}$ such that $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$. Thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{KPR}$, if $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}$.
(f) If $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$. If $\mathcal{M}$ and $\mathcal{N}$ are furthermore assumed to be models of $\mathrm{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, then (e) above implies that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{rank}}\mathcal{N}$.
(g) If both $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{KPR}$, and $\mathcal{M\preceq_{\Delta_{0}}N}$, then the statement “$\mathcal{N}$ is taller than $\mathcal{M}$” is equivalent to “there is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ that exceeds each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$”.
(h) A conservative extension is a faithful extension. The converse fails: Consider strongly inaccessible cardinals $\kappa$ and $\lambda$ with $\kappa<\lambda$, then $\left(\mathrm{V}_{\kappa},\in\right)\subsetneq_{\mathrm{faith}}\left(\mathrm{V}_{\lambda},\in\right)$ but by Tarski’s undefinability of truth $\left(\mathrm{V}_{\lambda},\in\right)$ is not a conservative extension of $\left(\mathrm{V}_{\kappa},\in\right)$ since the satisfaction predicate of $\left(\mathrm{V}_{\kappa},\in\right)$ is definable in $\left(\mathrm{V}_{\lambda},\in\right).$
(i) Generic extensions of models of ZF are end extensions but not rank extensions since they have the same ordinals. However, by (f) above, powerset preserving end extensions of models of KPR are rank extensions, thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{KPR}$ we have:

### v6

@@ 2. PRELIMINARIES
(d) We say that $X\subseteq M$ is coded (in $\mathcal{M)}$ if there is some $c\in M$ such that $\mathrm{Ext}_{\mathcal{M}}(c)=X.$
(e) Suppose $\mathcal{M}\models\mathsf{ZF}$. We say that $X\subseteq M$ is $\mathcal{M}$-amenable if $(\mathcal{M},X)\models\mathsf{ZF}(\mathrm{X})$; here $X$ is the interpretation of $\mathrm{X}$.
2.3. Definitions. (‘geometric shapes’ of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures, and $\mathcal{M}$ a submodel [^6] The notion of a submodel here is the usual one in model theory; i.e., $\mathcal{M}\subseteq\mathcal{N}$ means $M\subseteq N$ and the interpretation of each nonlogical symbol of $\mathcal{L}$ in $\mathcal{M}$ is the restriction to $M$ of the corresponding of interpretation in $\mathcal{N}$ (in particular, $\mathrm{Ext}_{\mathcal{M}}(a)\subseteq\mathrm{Ext}_{\mathcal{N}}(a)$ for all $a\in M).$ of $\mathcal{N}$ (written $\mathcal{M}\subseteq\mathcal{N)}.$
  1. (a)
    $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$ if $M^{\ast}=\bigcup\limits_{a\in M}\mathrm{Ext}_{\mathcal{N}}(a).$
  2. (b)
    Suppose $a\in M.$ $\mathcal{N}$ fixes $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{N}}(a)$, and $\mathcal{N}$ enlarges $a$ if $\mathrm{Ext}_{\mathcal{M}}(a)\subsetneq\mathrm{Ext}_{\mathcal{N}}(a).\vskip 3.0pt plus 1.0pt minus 1.0pt$
  3. (c)
    $\mathcal{N}$ end extends $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$), if $\mathcal{N}$ fixes every $a\in M$. End extensions are also referred to in the literature as transitive extensions, and in the old days as outer extensions.
  4. (d)
    $\mathcal{N}$ is a powerset-preserving end extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$) if $(i)$ $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $(ii)$ for if $a\in M$, $b\in N$, and $\mathcal{N}\models(b\subseteq a)$, then $b\in M$.
  5. (e)
@@
  6. (f)
    $\mathcal{N}$ is a cofinal extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{N)}$ if for every $b\in N$ there is some $a\in M$ such that $b\in\mathrm{Ext}_{\mathcal{N}}(a)$, i.e., $\mathcal{N}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$.
  7. (g)
    $\mathcal{N}$ is taller than $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{taller}}\mathcal{N)}$ if there is some $b\in M$ such that $M\subseteq\mathrm{Ext}_{\mathcal{N}}(b).$
2.4. Definitions. (‘logical shapes’ of extensions). Suppose $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, and $\mathcal{M}$ and $\mathcal{N}$ are $\mathcal{L}$-structures such that $\mathcal{M}$ is a submodel of $\mathcal{N}$.
  1. (a)
    For a subset $\Gamma$ of $\mathcal{L}$-formulae, $\mathcal{M}$ is a $\Gamma$-elementary submodel of $\mathcal{N}$ (written $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$) if for all $n$-ary formulae $\gamma\in\Gamma$ and for all $a_{1},\cdot\cdot\cdot,a_{n}$ in $M$, $\mathcal{M}\models\gamma(a_{1},\cdot\cdot\cdot,a_{n})$ iff $\mathcal{N}\models\gamma(a_{1},\cdot\cdot\cdot,a_{n}).$ $\mathcal{M}$ is an elementary submodel of $\mathcal{N}$ (written $\mathcal{M}\preceq\mathcal{N}$) if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ for $\Gamma=$ $\mathcal{L}$-formulae. We say that $\mathcal{M}$ is an elementary submodel of $\mathcal{N}$ if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ for the set $\Gamma$ of all $\mathcal{L}$-formulae (written $\mathcal{M}\preceq\mathcal{N}$). For a given $\gamma\in\Gamma$ we say that $\gamma$ is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\preceq_{\{\gamma\}}\mathcal{N}$. As usual, we write $\mathcal{M}\prec_{\Gamma}\mathcal{N}$ if $\mathcal{M}\preceq_{\Gamma}\mathcal{N}$ and $\mathcal{M}\subsetneq\mathcal{N}$ (similarly for $\mathcal{M}\prec\mathcal{N}$).
  2. (b)
    $\mathcal{N}$ is a self-extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}$) if $\mathcal{N}$ is interpretable in $\mathcal{M}$ (i.e., $\mathcal{N}$ is isomorphic to a model $\mathcal{N}^{\ast}$ whose universe $N^{\ast}$ of and the interpretation $R^{\mathcal{N}^{\ast}}$ of each $R\in\mathcal{L}$, are all $\mathcal{M}$-definable), and additionally, there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into $\mathcal{N}^{\ast}$. In this context, $\mathcal{N}$ is an elementary self-extension of $\mathcal{M}$ if the map $j$ is an elementary embedding.
  3. (c)
@@
    $\mathcal{N}$ is a conservative extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$) if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-definable.
  4. (d)
@@
    For $\mathcal{M}\models\mathsf{ZF}$, $\mathcal{N}$ is a faithful extension of $\mathcal{M}$ (written $\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N)}$, if for every $\mathcal{N}$-definable $D$, $M\cap D$ is $\mathcal{M}$-amenable (as in Definition 2.2(e)). [^8] The notion of a faithful extension was first introduced and studied in the context of elementary extensions in [En-3].
2.5. Remark. The following are readily verifiable:
  1. (a)
    If $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ is a chain of models extending $\mathcal{M}$ whose union is $\mathcal{N}$, and $\mathcal{M}\preceq_{\Pi_{k,\mathrm{cons}}}\mathcal{N}_{k}\preceq_{\Pi_{k}}\mathcal{N}_{k+1}$ for each $k\in\omega,$ then $\mathcal{M}\preceq_{\mathrm{cons}}\mathcal{N}$. [^9] This fact readily follows from Theorem 2.11. It comes handy in the first proof of Theorem 3.1. Here $\mathcal{M}\preceq_{\Pi_{k,\mathrm{cons}}}\mathcal{N}_{k}$ is shorthand for: $\mathcal{M}\preceq_{\Pi_{k}}\mathcal{N}_{k}$ and $\mathcal{M}\preceq_{{}_{\mathrm{cons}}}\mathcal{N}_{k}$.
  2. (b)
@@
    $\left(\mathcal{M}\subseteq_{\mathrm{self}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right).\vskip 3.0pt plus 1.0pt minus 1.0pt$
  3. (c)
    Assuming $\mathcal{M}\models\mathsf{ZF}$, $\left(\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}\right)\Rightarrow\left(\mathcal{M}\subseteq_{\mathrm{faith}}\mathcal{N}\right).$
The implications in (b) and (c) are not reversible. More specifically, (b) is not reversible since, e.g., if $\mathcal{N}$ is a self-extension of $\mathcal{M},$ then $\left|M\right|=\left|N\right|,$ but it is possible to build a conservative elementary extension $\mathcal{N}$ of a model $\mathcal{M}$ with $\left|M\right|<\left|N\right|,$ e.g., if $\mathcal{U}$ is an $\mathcal{M}$-ultrafilter, then the ultrapower formation modulo $\mathcal{U}$ can be iterated along any linear order (of arbitrary cardinality) and results in a conservative elementary extension of $\mathcal{M}$. In particular, it is possible to arrange $\mathcal{M}\subseteq_{\mathrm{cons}}\mathcal{N}$ such that $M$ is countable but $N$ is uncountable. Another way to see that $(i)$ is not reversible is to note that by a theorem of Gaifman [G-1] every elementary self-extension of a model $\mathcal{M}$ of the $\mathsf{ZF}$ is a cofinal extension. [^10] Indeed Gaifman’s theorem is stated for models $\mathcal{M}$ of the fragment $\mathsf{ZR}$ of $\mathsf{ZF}$, where $\mathsf{ZR}$ is the result of extending $\mathsf{Z}$ (Zermelo set theory) with an axiom that every element is a member of some set of the form $\mathrm{V}_{\alpha}.$ However, by Theorem 3.1 there are models of $\mathsf{ZF}$ that have taller conservative elementary extensions. To see that (c) is not reversible, note that if $\kappa$ and $\lambda$ are strongly inaccessible cardinals with $\kappa<\lambda$, then $\left(\mathrm{V}_{\kappa},\in\right)$ is faithfully extended by $\left(\mathrm{V}_{\lambda},\in\right)$, but since the truth predicate for $\left(\mathrm{V}_{\kappa},\in\right)$ is parametrically definable in $\left(\mathrm{V}_{\lambda},\in\right)$, by Tarski’s undefinability of truth $\left(\mathrm{V}_{\lambda},\in\right)$ is not a conservative extension of $\left(\mathrm{V}_{\kappa},\in\right)$.
It is also noteworthy that the most common examples of definable extensions in the context of models of set theory are elementary extensions. However, the Boolean-valued approach to forcing provides a wealth of definable extensions that are not elementary. More specifically, if $\mathcal{M}$ is a model of $\mathsf{ZFC}$, $\mathbb{B}$ is an element of $\mathcal{M}$ that is a complete Boolean algebra from the point of view of $\mathcal{M}$, and $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$ that is in $\mathcal{M}$, then there is an $\mathcal{M}$-definable embedding $j$ of $\mathcal{M}$ into the model $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ (where $\mathcal{M}^{\mathbb{B}}$ is the $\mathbb{B}$-valued forcing extension of the universe, as calculated in $\mathcal{M}$). However, in typical cases $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ are not elementarily equivalent, e.g., we can always arrange $\mathbb{B}$ so that the truth-value of the continuum hypothesis changes in the transition between the $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$. Furthermore, even though $\mathcal{M}$ and $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$ might be elementarily equivalent in some cases, in general the embedding $j$ is not elementary if $\mathbb{B}$ is a nontrivial notion of forcing, since the statement “there is a filter over $\mathbb{B}$ that meets all the dense subsets of $\mathbb{B}$” holds in $\mathcal{M}^{\mathbb{B}}/\mathcal{U}$, but not in $\mathcal{M}$ (assuming $\mathbb{B}$ is a nontrivial notion of forcing).
  1. (a)
    For each $n\in\omega,$ $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{N}$ iff $\mathcal{M}\preceq_{\Pi_{n}}\mathcal{N}.\vskip 3.0pt plus 1.0pt minus 1.0pt$
  2. (b)
    $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, implies $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$; and $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$ implies $\mathcal{M}\preceq_{\Delta_{1}}\mathcal{N}$.
  3. (c)
    Suppose $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, $\mathcal{N}$ fixes $a\in M$, and there is bijection $f$ in $\mathcal{M}$ between $a$ and some $b\in M$. Then $\mathcal{N}$ fixes $b$ as well (by $\Delta_{0}$-elementarity, $f$ remains a bijection between $a$ and $b$ as viewed from $\mathcal{N}$).
  4. (d)
    Suppose $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{KPR}$, $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, and $\alpha\in\mathrm{Ord}^{M}$. Then $\mathrm{V}_{\alpha}^{\mathcal{M}}=\mathrm{V}_{\alpha}^{\mathcal{N}}$; this follows from the fact that the formula expressing that $x$ is an ordinal and $y=\mathrm{V}_{x}$ as a $\Sigma_{1}^{\mathcal{P}}$-formula in $\mathsf{KPR}$ (using the usual recursive definition of the $\mathrm{V}_{\alpha}$-hierarchy).
  5. (e)
    If $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, then $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$. On the other hand, since the formula expressing “$x$ is an ordinal and $y=\mathrm{V}_{x}$” can be written as a $\Sigma_{1}^{\mathcal{P}}$-formula, it is absolute between models $\mathcal{M}$ and $\mathcal{N}$ such that $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$. Thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$:
    $\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}\Rightarrow\mathcal{M}\subseteq_{\mathrm{rank}}\mathcal{N}.$
    As pointed out by the referee, by [M-1, Proposition 6.17], the converse of the above implication also holds provided $\mathcal{M}\models\mathsf{Most}$ and $\mathcal{N}\models\mathsf{Mac}$.
  6. (f)
    If $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$, then $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$. If $\mathcal{M}$ and $\mathcal{N}$ are furthermore assumed to be models of $\mathsf{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$, then (e) above implies that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{rank}}\mathcal{N}$.
  7. (g)
    If both $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{KPR}$, and $\mathcal{M}\preceq_{\Delta_{0}}\mathcal{N}$, then the statement $\mathcal{N}$ is taller than $\mathcal{M}$” is equivalent to “there is some $\gamma\in\mathrm{Ord}^{\mathcal{N}}$ that exceeds each $\alpha\in\mathrm{Ord}^{\mathcal{M}}$”.
  8. (h)
    Generic extensions of models of $\mathsf{ZF}$ are end extensions but not rank extensions since they have the same ordinals. However, by (f) above, powerset preserving end extensions of models of $\mathsf{KPR}$ are rank extensions, thus for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$ we have:
@@
    On the other hand, since $\mathcal{P}(x)=y$ is $\Pi_{1}$-statement, we have:
    $\mathcal{M}\preceq_{\Sigma_{1},\mathrm{end}}\mathcal{N}\Longrightarrow\mathcal{M}\subseteq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}.$
    The converse of the above implication need not be true for arbitrary models $\mathcal{M}$ and $\mathcal{N}$ of $\mathsf{KPR}$. The referee has offered the following counterexample: let $\mathcal{N}$ be a model of $\mathsf{ZFC}$ that is $\omega$-standard but whose $\omega_{1}$ is nonstandard, and let $\mathcal{M}$ be the well-founded part of $\mathcal{N}$. Then $\mathcal{M}\models\mathsf{KPR}$, and the property of being equinumerous to a von Neumann ordinal is a $\Sigma_{1}$-property of $\mathcal{P}^{\mathcal{N}}(\omega)$ that holds in $\mathcal{N}$ but fails in $\mathcal{M}$. However, if $\mathcal{M}\models\mathsf{Most}$ and $\mathcal{N}\models\mathsf{Mac}$, then by [M-1, Proposition 6.17] the converse of the above implication holds.
  9. (i)
    As we shall see in Lemma 4.3, faithful end extensions of models of $\mathsf{ZF}$ are rank extensions.
  10. (g)
    Thanks to the provability of the induction scheme (over natural numbers) in $\mathsf{ZF}$, if $\mathcal{N}$ is a faithful extension of a model $\mathcal{N}$ (where $\mathcal{M}$ is a model of $\mathsf{ZF}$), then $\mathcal{N}$ fixes each element of $\omega^{\mathcal{M}}$ (but $\mathcal{N}$ need not fix $\omega^{\mathcal{M}}$, e.g., consider the case when $\mathcal{N}$ is an internal ultrapower of $\mathcal{M}$ modulo a nonprincipal ultrafilter over $\omega^{\mathcal{M}}$).
2.7. Definition. Reasoning within $\mathsf{KP}$, for each object $a$ in the universe of sets, let $\dot{a}$ be a constant symbol denoting $a$ (where the map $a\mapsto\dot{a}$ is $\Delta_{1}$, e.g., $\dot{a}=\left\langle 3,a\right\rangle$ as in Devlin’s monograph [D]). Let $\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(x)$ be the $\mathcal{L}_{\mathrm{set}}$-formula that defines the proper class $\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}$ of sentences in the language
$\mathcal{L}^{+}=$ $\mathcal{L}\cup\{\dot{a}:a\in\mathrm{V}\}$,